Find the volume of a regular hexagonal pyramid with base edge 8cm and height 12cm.
Answer by mananth(16946)


V = 1/3 * b*h
B= area of base
h = height of prism
Area of base = 3*Sqrt(3)/2 * a^2
3*sqrt(3)/2 * 8^2
96 sqrt(3)
V = 1/3 * b * h
V = 1/3 * 96*sqrt(3)* 12
= 665.107 cm^3